# Topology of Reasoning: Understanding Large Reasoning Models through Reasoning Graph Properties

### v2

@@ Abstract
Recent large-scale reasoning models have achieved state-of-the-art performance on challenging mathematical benchmarks, yet the internal mechanisms underlying their success remain poorly understood. In this work, we introduce the notion of a reasoning graph, extracted by clustering hidden-state representations at each reasoning step, and systematically analyze three key graph-theoretic properties: cyclicity, diameter, and small-world index, across multiple tasks (GSM8K, MATH500, AIME 2024). Our findings reveal that distilled reasoning models (e.g., DeepSeek-R1-Distill-Qwen-32B) exhibit significantly more recurrent cycles (about 5 per sample), substantially larger graph diameters, and pronounced small-world characteristics (about 6x) compared to their base counterparts. Notably, these structural advantages grow with task difficulty and model capacity, with cycle detection peaking at the 14B scale and exploration diameter maximized in the 32B variant, correlating positively with accuracy. Furthermore, we show that supervised fine-tuning on an improved dataset systematically expands reasoning graph diameters in tandem with performance gains, offering concrete guidelines for dataset design aimed at boosting reasoning capabilities. By bridging theoretical insights into reasoning graph structures with practical recommendations for data construction, our work advances both the interpretability and the efficacy of large reasoning models. Implementation available here: https://github.com/gouki510/Topology_of_Reasoning
@@ 1 Introduction
First, we extract reasoning graph nodes by clustering hidden states of LLMs using $K$-means during reasoning tasks. Then, for each reasoning task sample, we construct a reasoning graph by connecting the nodes visited by the model during inference and analyze its properties. Visualizing the reasoning graph with t-SNE clearly demonstrates that reasoning graphs from large reasoning models include cycles and have a broader exploration range compared to base models. Quantitatively assessing these cyclic properties of reasoning graph confirms that large reasoning models exhibit significantly more cycles than base models. Furthermore, the proportion of reasoning graphs containing cycles increases progressively with task difficulty, as observed across datasets GSM8K [6] (easier), MATH500 [19] (intermediate), and AIME 2024 [10] (more challenging). To quantitatively examine whether large reasoning models explore a broader range of nodes during inference, we compared the diameters of their reasoning graphs. We observed notably larger diameters for large reasoning models, suggesting that they explore a wider variety of reasoning states, potentially enabling more sophisticated inference strategies. To gain deeper insights into the structural underpinnings of reasoning ability, we assess the small-world characteristics of these graphs, revealing that large reasoning models construct graphs exhibit significantly higher small-world properties. This small-world structures indicate that reasoning graph of large reasoning model have dense local clustering structures which likely contribute to improved reasoning performance. These graph-theoretic properties become more pronounced with increasing model size, suggesting that greater model capacity facilitates the formation of graph structures beneficial for reasoning. These distinct graph-theoretic characteristics provide critical insights into the mechanisms underlying enhanced reasoning performance.
Furthermore, to connect insights derived from reasoning graph properties directly to practical LLM training, we examine reasoning graphs through the lens of FT, particularly analyzing the s1 method [35]. Training the base model using the s1 dataset leads to increased graph diameters, and notably, training on the improved s1-v1.1 dataset, which achieves higher performance, results in even larger graph diameters. These results indicate that more effective SFT data for reasoning can be characterized by distinctive graph-theoretic properties, suggesting practical guidance for designing better data construction methods.
@@
  • Graph-Theoretic Analysis of Enhanced Reasoning (Section 4): Through empirical analyses across multiple mathematical reasoning datasets, we identify distinctive graph-theoretic properties of large reasoning models, including (1) increased cyclic behavior, (2) larger graph diameters, and (3) heightened small-world characteristics. These structural patterns offer key insights into the mechanisms behind recent breakthroughs in the reasoning performance of LLMs, highlighting how advanced models explore a broader range of reasoning states and transition between them more effectively.
  • Insights for Effective SFT Data Construction (Section 5): By examining how SFT datasets influence reasoning graph characteristics, we reveal that refined datasets lead to larger graph diameters and better reasoning performance. These insights provide actionable guidelines for designing training data to explicitly enhance reasoning capabilities.
@@ 2 Related Works
2.1 Approaches to Enhanced Reasoning
Recent methods for enhancing reasoning in language models include: (1) search-based methods that utilize additional inference-time computation or iterative self-improvement, and (2) reinforcement learning (RL)-based fine-tuning, which has notably driven significant performance breakthroughs.
Search-based methods involve external computation of inference time such as parallel sampling [29, 4], sophisticated verifier-based searches [32, 54, 46], or internal in-context refinement and self-correction strategies [15, 16, 55, 28, 22]. Despite their effectiveness, these methods often require careful design or redundant computation. RL methods autonomously discover effective reasoning strategies, encompassing both off-policy [58, 20] and on-policy techniques [59, 25, 8]. Recent advances like DeepSeek-R1 [9], leveraging algorithms such as PPO [39] and GRPO [40], demonstrate significant improvements through structured reasoning traces [30, 57]. A notable phenomenon is the emergence of the “aha moment”, characterized by models spontaneously revising their reasoning strategies, inspiring new training approaches such as s1 [35] and Think-DPO [53].
@@ 2.1 Approaches to Enhanced Reasoning
2.2 Analytical Studies on Reasoning Capabilities
Earlier studies prior to recent breakthroughs in RL-based reasoning emphasized reasoning graphs to explain LLM capabilities. Previous work [37] theoretically and empirically demonstrated that reasoning capabilities emerge due to the locality property inherent in natural language data; specifically, models achieve better accuracy by traversing intermediate variables frequently co-occurring during training. Additionally, Wang et al. [47] proposed extracting reasoning graphs by clustering internal model states using $K$-means, hypothesizing that pre-training data enables a random walk over these graph nodes. Other studies [12, 51, 5, 45] have created simple toy tasks with explicit reasoning graphs to better understand the mechanisms underlying reasoning abilities in language models.
More recently, DeepSeek-R1 [9] have demonstrated significant improvements through RL without explicit reasoning supervision. Analyses following this advancement have explored steering vectors [44], cognitive behaviors [17], and anthropomorphic expressions [52]. However, while recent reasoning models are featured with enhanced reasoning traces, no studies have analyzed them through reasoning graph perspective. Addressing this gap is essential for understanding current advancements.
@@ 3.1 Extracting Reasoning Graph from LLM’s Representations
Following the previous work [47], we aggregate all segment representations $\mathcal{S}=\bigl{\{}\,s_{i}^{\,\ell}\mid 1\leq i\leq T,\;x\in\mathcal{D}\bigr{\}}$ and run $K$-means (default $K=200$) to obtain clusters $\{C_{k}\}_{k=1}^{K}$ with centroids $\{c_{k}\}$. Each centroid $c_{k}$ corresponds to a node $v_{k}$ in the reasoning graph:
@@
The distance $d(v_{i},v_{j})$ between nodes $v_{i}$ and $v_{j}$ is defined as the Euclidean distance between their corresponding centroids $c_{i}$ and $c_{j}$. Figure 2-(b) presents representative nodes obtained from clustering the DeepSeek-R1-Distill-Qwen-32B [9] using GSM8K dataset [6]. Each clustered node corresponds to simple computations encountered within tasks. As characteristic of reasoning models, some nodes include the term “wait”, indicative of an “aha moment” where the model rechecks its outputs.
@@
Informally, the edges in the reasoning graph represent the sequential path of nodes visited by the model for each question during inference. Formally, for each question $x$, let $\pi=(i_{1},i_{2},\dots,i_{T}),$ where $\,i_{j}=\arg\min_{k}\|s_{j}^{\,\ell}-c_{k}\|_{2}$ assigns segment $r_{j}$ to its nearest centroid. We then define the directed-edge set
$E=\bigl{\{}(v_{i_{j}}\to v_{i_{j+1}})\mid j=1,\dots,T-1\bigr{\}},$
@@ 3.2 Measuring Graph Properties
Having extracted reasoning graphs from LLM representations, we evaluate their structural properties from three perspectives: (1) Cycles, (2) Diameter, and (3) Small-World index. Simple implementations of each method are provided in Appendix B.
@@
where $d(u,v)$ is the shortest-path distance from node $u$ to node $v$. Letting $N$ be the total number of nodes and $K$ the mean degree of the undirected graph, we approximate the corresponding Erdős–Rényi random-graph baseline values [3] as $C_{\mathrm{rand}}=\frac{K}{N-1},L_{\mathrm{rand}}=\frac{\ln N}{\ln K},$ and define the small-world index by $S=\frac{C/C_{\mathrm{rand}}}{L/L_{\mathrm{rand}}}.$ The clustering coefficient describes the tendency of nodes to form tightly interconnected groups, while the average path length indicates how efficiently information propagates through the network. The small-world index combines these characteristics, highlighting a graph’s ability to maintain local cohesion while supporting rapid global connectivity.
@@ 4 Analyzing Enhanced Reasoning through Graph-Theoretic Properties
Qwen32B
sample idx=5
sample idx=16
sample idx=20
sample idx=30
sample idx=34
@@
sample idx=5
sample idx=16
sample idx=20
sample idx=30
sample idx=34
@@
We utilize the Qwen2.5 family distilled from DeepSeek-R1 [9] as our large reasoning models, available in sizes of 1.5B, 7B, 14B, and 32B parameters. The corresponding base models and their details are provided in Appendix C. By default, we use the highest-performing 32B variant. Unless otherwise specified, we extract reasoning graphs from the hidden layer positioned at 90% depth (e.g., layer 58 in the 64-layer 32B model). We employ the GSM8K [6], MATH500 [19], and AIME 2024 [10] datasets for constructing the reasoning graphs.
@@ 4.1 Visualization and Quantification of Cycles in Reasoning Graphs
To quantitatively validate these qualitative observations, we employed the cycle detection method introduced in Section 3.2. Figure 4 shows cycle detection rates for GSM8K, MATH500, and AIME 2024 datasets, comparing the large reasoning model (blue) with the base model (orange). The horizontal axis denotes different relative depths of hidden layers (0.1, 0.3, 0.5, 0.7, and 0.9), corresponding respectively to layers 6, 19, 32, 45, and 58 in the 64-layer Qwen2.5-32B. Across all layers, the large reasoning model consistently exhibited a notably higher frequency of cyclic reasoning graphs compared to the base model. Additionally, we observed higher cycle detection rates at the earlier and later layers, with lower detection rates in intermediate layers. This pattern suggests that intermediate layers compress token representations, making less cycle detection difficult, whereas layers closer to input or output exhibit clearer cyclic behaviors. The results for varying the hyperparameter $k$ of the $K$-means clustering are provided in Appendix D, showing consistent trends across all tested values of $k$. Furthermore, another consistent trend emerges in which cycle detection ratios increase with the increasing complexity of tasks, progressing from GSM8K through MATH500 to AIME 2024. These findings reinforce the hypothesis that cycles within reasoning graphs contribute to the enhanced reasoning capabilities observed in large reasoning models.
@@ 4.2 Analyzing Reasoning Exploration through Graph Diameter
Moreover, we observed a clear trend of progressively increasing graph diameters in deeper hidden layers, suggesting that richer contextual representations at deeper layers correspond to broader exploration scopes. These observations imply similarities between expanding the number of output tokens (thus enlarging the exploration scope) and increasing model depth from the perspective of reasoning graph diameters. Our findings suggest a unified explanation based on reasoning graph diameters, which aligns closely with recent studies that emphasize improved reasoning through iterative deep-layer processing [33, 18]. Additional results for the MATH500 and AIME 2024 datasets can be found in Appendix E, demonstrating the same trends across all tasks.
@@ 4.3 Small-World Structure and Enhanced Reasoning Efficiency
Figure 6-(a) depicts the distributions of clustering coefficients and average path lengths for the large reasoning model and the base model at a layer ratio of 0.9. The large reasoning model clearly demonstrates notably higher clustering coefficients alongside longer average path lengths. This combination indicates that reasoning graphs in large reasoning models form densely interconnected local clusters while also having some nodes connected by relatively long-range paths. Such a structure allows quick access via short paths to arbitrary nodes within local clusters, facilitating easier recovery from incorrect reasoning pathways and potentially enhancing reasoning performance. This structural pattern aligns with recent theoretical findings [26] that model reasoning processes as Markov chains, comprising densely connected nodes (representing simple reasoning steps) and sparsely connected critical transitions (representing complex reasoning steps). Additional results for other layer ratios are provided in Appendix F, consistently showing similar trends across all layers.
@@ 4.4 Impact of Model Size on Reasoning Graph Properties
Figure 7: (a) Relationship between cycle detection ratio and AIME 2024 accuracy across different model sizes. The cycle detection ratio generally increases with model size up to 14B, which achieves a 100% cycle detection ratio. However, the 32B model, which achieves the highest accuracy, shows a slightly lower cycle detection ratio compared to the 14B model. (b) Relationship between cycle count and accuracy across different model sizes. Larger models demonstrate increased cycle counts, with the 32B model, which achieves the highest accuracy, exhibiting the greatest number of cycles. (c) Relationship between reasoning graph diameter and accuracy across different model sizes. The 32B model, achieving the highest accuracy, also exhibits the largest graph diameter.
To clarify how model size influences reasoning graph properties, we analyzed the relationship between cycle detection ratios, reasoning graph diameters, cycle counts, and task accuracy on the AIME 2024 dataset. The results of the small-world index relative to model size are detailed in Appendix G.

### v3

@@ Abstract
Recent large-scale reasoning models have achieved state-of-the-art performance on challenging mathematical benchmarks, yet the internal mechanisms underlying their success remain poorly understood. In this work, we introduce the notion of a reasoning graph, extracted by clustering hidden‐state representations at each reasoning step, and systematically analyze three key graph-theoretic properties: cyclicity, diameter, and small-world index, across multiple tasks (GSM8K, MATH500, AIME 2024). Our findings reveal that distilled reasoning models (e.g., DeepSeek-R1-Distill-Qwen-32B) exhibit significantly more recurrent cycles (about 5 per sample), substantially larger graph diameters, and pronounced small-world characteristics (about 6x) compared to their base counterparts. Notably, these structural advantages grow with task difficulty and model capacity, with cycle detection peaking at the 14B scale and exploration diameter maximized in the 32B variant, correlating positively with accuracy. Furthermore, we show that supervised fine-tuning on an improved dataset systematically expands reasoning graph diameters in tandem with performance gains, offering concrete guidelines for dataset design aimed at boosting reasoning capabilities. By bridging theoretical insights into reasoning graph structures with practical recommendations for data construction, our work advances both the interpretability and the efficacy of large reasoning models. Implementation available here: https://github.com/gouki510/Topology_of_Reasoning
@@ 1 Introduction
First, we extract reasoning graph nodes by clustering hidden states of LLMs using kmeans during reasoning tasks. Then, for each reasoning task sample, we construct a reasoning graph by connecting the nodes visited by the model during inference and analyze its properties. Visualizing the reasoning graph with t-SNE clearly demonstrates that reasoning graphs from large reasoning models include cycles and have a broader exploration range compared to base models. Quantitatively assessing these cyclic properties of reasoning graph confirms that large reasoning models exhibit significantly more cycles than base models. Furthermore, the proportion of reasoning graphs containing cycles increases progressively with task difficulty, as observed across datasets GSM8K [7] (easier), MATH500 [23] (intermediate), and AIME 2024 [11] (more challenging). To quantitatively examine whether large reasoning models explore a broader range of nodes during inference, we compared the diameters of their reasoning graphs. We observed notably larger diameters for large reasoning models, suggesting that they explore a wider variety of reasoning states, potentially enabling more sophisticated inference strategies. To gain deeper insights into the structural underpinnings of reasoning ability, we assess the small-world characteristics of these graphs, revealing that large reasoning models construct graphs exhibit significantly higher small-world properties. This small-world structures indicate that reasoning graph of large reasoning model have dense local clustering structures which likely contribute to improved reasoning performance. These graph-theoretic properties become more pronounced with increasing model size, suggesting that greater model capacity facilitates the formation of graph structures beneficial for reasoning. These distinct graph-theoretic characteristics provide critical insights into the mechanisms underlying enhanced reasoning performance.
Furthermore, to connect insights derived from reasoning graph properties directly to practical LLM training, we examine reasoning graphs through the lens of FT, particularly analyzing the s1 method [44]. Training the base model using the s1 dataset leads to increased graph diameters, and notably, training on the improved s1-v1.1 dataset, which achieves higher performance, results in even larger graph diameters. These results indicate that more effective SFT data for reasoning can be characterized by distinctive graph properties, suggesting practical guidance for designing better data construction methods.
@@
  • Graph-Theoretic Analysis of Enhanced Reasoning (Section˜4): Through empirical analyses across multiple datasets, we identify distinctive graph properties of large reasoning models, including (1) increased cyclic behavior, (2) larger graph diameters, and (3) heightened small-world characteristics. These structural patterns offer key insights into the mechanisms behind recent breakthroughs in the reasoning performance of LLMs, highlighting how advanced models explore a broader range of reasoning states and transition between them more effectively.
  • Insights for Effective SFT Data Construction (Section˜5): By examining how SFT datasets influence reasoning graph characteristics, we reveal that refined datasets lead to larger graph diameters and better reasoning performance. These insights provide actionable guidelines for designing training data to explicitly enhance reasoning capabilities.
@@ 2 Related Works
Approaches to Enhanced Reasoning
Recent methods for enhancing reasoning in language models include: (1) search-based methods that utilize additional inference-time computation or iterative self-improvement, and (2) reinforcement learning (RL)-based fine-tuning, which has notably driven significant performance breakthroughs. Search-based methods involve external computation of inference time such as parallel sampling [34, 5], sophisticated verifier-based searches [38, 70, 62], or internal in-context refinement and self-correction strategies [17, 18, 71, 33, 27]. Despite their effectiveness, these methods often require careful design or redundant computation. RL methods autonomously discover effective reasoning strategies, encompassing both off-policy [74, 24] and on-policy techniques [75, 30, 9]. Recent advances like DeepSeek-R1 [10], leveraging algorithms such as PPO [50] and GRPO [51], demonstrate significant improvements through structured reasoning traces [35, 73]. A notable phenomenon is the emergence of the “aha moment”, characterized by models spontaneously revising their reasoning strategies, inspiring new training approaches such as s1 [44] and Think-DPO [69].
@@
Analytical Studies on Reasoning Capabilities
Earlier studies prior to recent breakthroughs in RL-based reasoning emphasized reasoning graphs to explain LLM capabilities. Previous work [48] theoretically and empirically demonstrated that reasoning capabilities emerge due to the locality property inherent in natural language data; specifically, models achieve better accuracy by traversing intermediate variables frequently co-occurring during training. Additionally, Wang et al. [63] proposed extracting reasoning graphs by clustering internal model states using $K$-means, hypothesizing that pre-training data enables a random walk over these graph nodes. Other studies [13, 67, 6, 59] have created simple toy tasks with explicit reasoning graphs to better understand the mechanisms underlying reasoning abilities in language models. More recently, DeepSeek-R1 [10] have demonstrated significant improvements through RL without explicit reasoning supervision. Analyses following this advancement have explored steering vectors [58], cognitive behaviors [19], and anthropomorphic expressions [68].
However, while recent reasoning models are featured with enhanced reasoning traces, no studies have analyzed them through reasoning graph perspective. Addressing this gap is essential for understanding current advancements.
@@ 3.1 Extracting Reasoning Graph from LLM’s Representations
Following the previous work [63], we aggregate all segment representations $\mathcal{S}=\bigl\{\,s_{t}^{\,\ell}\mid 1\leq t\leq T,\;x\in\mathcal{D}\bigr\}$ and run $K$-means (default $K=200$) to obtain clusters $\{C_{k}\}_{k=1}^{K}$ with centroids $\{c_{k}\}$. Each centroid $c_{k}$ corresponds to a node $v_{k}$ in the reasoning graph:
@@
The distance $d(v_{i},v_{j})$ between nodes $v_{i}$ and $v_{j}$ is defined as the Euclidean distance between their corresponding centroids $c_{i}$ and $c_{j}$. Figure 2-(b) presents representative nodes obtained from clustering the DeepSeek-R1-Distill-Qwen-32B [10] using GSM8K dataset [7]. Each clustered node corresponds to simple computations encountered within tasks. As characteristic of reasoning models, some nodes include the term “wait”, indicative of an “aha moment” where the model rechecks its outputs. A more fine-grained analysis, leveraging an LLM-as-judge to characterize the semantics of these clusters, is provided in Appendix B.
@@
Informally, the edges in the reasoning graph represent the sequential path of nodes visited by the model for each question during inference. Formally, for each question $x$, let $\pi=(i_{1},i_{2},\dots,i_{T}),$ where $i_{t}=\arg\min_{k}\|s_{t}^{\,\ell}-c_{k}\|_{2}$ assigns segment $r_{t}$ to its nearest centroid. We then define the directed-edge set
$E=\bigl\{(v_{i_{t}}\to v_{i_{t+1}})\mid t=1,\dots,T-1\bigr\},$
@@ 3.2 Measuring Graph Properties
Having extracted reasoning graphs from LLM representations, we evaluate their structural properties from three perspectives: (1) Cycles, (2) Diameter, and (3) Small-World index. Simple implementations of each method are provided in Appendix C.
@@
where $d(u,v)$ is the shortest-path distance from node $u$ to node $v$. Letting $N$ be the total number of nodes and $K$ the mean degree of the undirected graph, we approximate the corresponding Erdős–Rényi random‐graph baseline values [4] as $C_{\mathrm{rand}}=\frac{K}{N-1},L_{\mathrm{rand}}=\frac{\ln N}{\ln K},$ and define the small‐world index by $S=\frac{C/C_{\mathrm{rand}}}{L/L_{\mathrm{rand}}}.$ The clustering coefficient describes the tendency of nodes to form tightly interconnected groups, while the average path length indicates how efficiently information propagates through the network. The small-world index combines these characteristics, highlighting a graph’s ability to maintain local cohesion while supporting rapid global connectivity.
@@ 4 Analyzing Enhanced Reasoning through Graph-Theoretic Properties
Qwen2.5-32B
sample #5
sample #21
sample #45
sample #211
sample #689
@@
sample #5
sample #21
sample #45
sample #211
sample #689
@@
We utilize the Qwen2.5 family distilled from DeepSeek-R1 [10] as our large reasoning models, available in sizes of 1.5B, 7B, 14B, and 32B parameters. The corresponding base models and their details are provided in Appendix D. By default, we use the highest-performing 32B variant. Unless otherwise specified, we extract reasoning graphs from the hidden layer positioned at 90% depth (e.g., layer 58 in the 64-layer 32B model). We employ the GSM8K [7], MATH500 [23], and AIME 2024 [11] datasets for constructing the reasoning graphs. Additional analyses on non-mathematical tasks, including StrategyQA [22] and LogicalDeduction from BIG-Bench [54], are provided in Appendix E.
@@ 4.1 Visualization and Quantification of Cycles in Reasoning Graphs
To quantitatively validate these qualitative observations, we employed the cycle detection method introduced in Section˜3.2. Figure 4 shows cycle detection rates for GSM8K, MATH500, and AIME 2024 datasets, comparing the large reasoning model (blue) with the base model (orange). The horizontal axis denotes different relative depths of hidden layers (0.1, 0.3, 0.5, 0.7, and 0.9), corresponding respectively to layers 6, 19, 32, 45, and 58 in the 64-layer Qwen2.5-32B. Across all layers, the large reasoning model consistently exhibited a notably higher frequency of cyclic reasoning graphs compared to the base model. Additionally, we observed higher cycle detection rates at the earlier and later layers, with lower detection rates in intermediate layers. This pattern suggests that intermediate layers compress token representations, making less cycle detection difficult, whereas layers closer to input or output exhibit clearer cyclic behaviors. The results for varying the hyperparameter $k$ of the $K$-means clustering are provided in Appendix F, showing consistent trends across all tested values of $k$. Furthermore, another consistent trend emerges in which cycle detection ratios increase with the increasing complexity of tasks, progressing from GSM8K through MATH500 to AIME 2024. These findings reinforce the hypothesis that cycles within reasoning graphs contribute to the enhanced reasoning capabilities observed in large reasoning models.
@@ 4.2 Analyzing Reasoning Exploration through Graph Diameter
Moreover, we observed a clear trend of progressively increasing graph diameters in deeper hidden layers, suggesting that richer contextual representations at deeper layers correspond to broader exploration scopes. These observations imply similarities between expanding the number of output tokens (thus enlarging the exploration scope) and increasing model depth from the perspective of reasoning graph diameters. Our findings suggest a unified explanation based on reasoning graph diameters, which aligns closely with recent studies that emphasize improved reasoning through iterative deep-layer processing [42, 21]. Additional results for the MATH500 and AIME 2024 datasets can be found in Appendix G, demonstrating the same trends across all tasks.
@@ 4.3 Small-World Structure and Enhanced Reasoning Efficiency
Figure 6-(a) depicts the distributions of clustering coefficients and average path lengths for the large reasoning model and the base model at a layer ratio of 0.9. The large reasoning model clearly demonstrates notably higher clustering coefficients alongside longer average path lengths. This combination indicates that reasoning graphs in large reasoning models form densely interconnected local clusters while also having some nodes connected by relatively long-range paths. Such a structure allows quick access via short paths to arbitrary nodes within local clusters, facilitating easier recovery from incorrect reasoning pathways and potentially enhancing reasoning performance. This structural pattern aligns with recent theoretical findings [31] that model reasoning processes as Markov chains, comprising densely connected nodes (representing simple reasoning steps) and sparsely connected critical transitions (representing complex reasoning steps). Additional results for other layer ratios are provided in Appendix H, consistently showing similar trends across all layers.
@@ 4.4 Impact of Model Size on Reasoning Graph Properties
Figure 7: (a) Relationship between cycle detection ratio and AIME 2024 accuracy across different model sizes. The cycle detection ratio generally increases with model size up to 14B, which achieves a 100% cycle detection ratio. (b) Relationship between cycle count and accuracy across different model sizes. Larger models demonstrate increased cycle counts, with the 32B model, which achieves the highest accuracy, exhibiting the greatest number of cycles. (c) Relationship between reasoning graph diameter and accuracy across different model sizes. The 32B model, achieving the highest accuracy, also exhibits the largest graph diameter.
To clarify how model size influences reasoning graph properties, we analyzed the relationship between cycle detection ratios, reasoning graph diameters, cycle counts, and task accuracy on the AIME 2024 dataset. The results of the small-world index relative to model size are detailed in Appendix I.